# Nonlinearly Preconditioned Gradient Methods under Generalized Smoothness

### v1

@@ 3.2 Convex setting
In the convex setting, Proposition 2.2 establishes a useful connection between anisotropic smoothness and the convexity of $f^{*}-L^{-1}(\bar{L}\star\phi^{*})$, i.e. the strong convexity of $f^{*}$ relative to $\bar{L}\star\phi^{*}$ with constant $L^{-1}$ in the Bregman sense. We utilize this connection in Proposition 3.5 and Theorem 3.6 in order to obtain sublinear convergence rates in the isotropic case.
@@
Obtaining standard sublinear rates for the function values is not straightforward. To the best of our knowledge, there do not exist such guarantees for the full generality of the setting we consider in this paper. Proposition 3.5 is useful in that regard, since it allows us to show a $O(1/k)$ rate for the suboptimality gap in the isotropic case.
@@
Let $f$ be convex and $\phi=h\circ\|\cdot\|$ with $h$ satisfying Assumption 1.1. For $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ the sequence of iterates generated by (2) with $\gamma=L^{-1}$ and $\lambda=\bar{L}^{-1}$, the following holds:
@@
The result above strengthens the known results regarding the convergence of the method from [maddison2021dual] and [laude2022anisotropic], while also answering the question posed in [maddison2021dual, p. 17]. We remark that although the obtained rate in (12) depends on the initial norm of the gradient, one can use the techniques from [vankov2024optimizing] or [gorbunov2024methods] to achieve a better complexity when specifying the reference function $\phi$. Nevertheless, such an endeavor is beyond the scope of this paper.
Although Theorem 3.6 provides a sublinear rate for the isotropic case, obtaining such guarantees for separable reference functions is not straightforward: key in the proof of Theorem 3.6 is the fact that $\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))={h^{*}}^{\prime}(\|\bar{L}^{-1}\nabla f(x^{k})\|)\overline{\operatorname{sgn}}(\nabla f(x^{k}))$ and therefore the convex gradient inequality for $f$ can directly be utilized. Nevertheless, we are able to prove the sublinear convergence rate for the suboptimality gap for subhomogeneous [aze1995uniformly, p. 708] reference functions using a different technique based on generalized conjugacy. More precisely, we utilize the characterization of anisotropically smooth functions as (generalized) envelopes and interpret the algorithm as a nonlinear proximal point method. Then, we combine the subhomogeneity of $\phi$ with the proof technique of [doikov2020inexact] for the inexact tensor methods and obtain the claimed rate. This result is captured in the following theorem.
@@
Let $f$ be convex and $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ be the sequence of iterates generated by (2) with $\gamma=L^{-1}$, $\lambda=\bar{L}^{-1}$ and assume that $\operatorname{dom}\phi=\mathbb{R}^{n}$. Moreover, let $\phi$ be $2$-subhomogeneous, i.e. such that $\phi(\theta x)\leq\theta^{2}\phi(x)$ for all $\theta\in[0,1]$. Then, for all $K\geq 1$
@@
where $\mathcal{D}_{0}=\sup\{\bar{L}(L^{-1}\star\phi)(x-x^{\star}):f(x)\leq f(x^{0})\}$.
In general the set $\mathcal{D}_{0}$ might be unbounded, except if $f$ has bounded level-sets, which is the case if $\operatorname*{arg\,min}f$ is bounded. In theory, this dependence on the initial level-set can be eliminated by considering an averaging procedure akin to [doikov2020inexact, Algorithm 3], thus leading to a convergence rate in terms of some function of the initial distance to the solution. Nevertheless, such methods tend to underperform in practice compared to their more straightforward counterparts.
@@
The function $h(x):=\cosh(x)-1$ is $2$-subhomogeneous, i.e. the following inequality holds:
@@ 4.1 Norm to power
For the first part of our experiments we consider the toy example of minimizing $f(x)=\tfrac{1}{4}\|x\|^{4}$, with $x\in\mathbb{R}^{500}$, using different preconditioning schemes. We consider the reference functions $\cosh(\|x\|)-1$, $\exp(\|x\|)-\|x\|-1$ and $-\|x\|-\ln(1-\|x\|)$ and remind that the algorithm generated by the latter function is a tighter version of the algorithms proposed in [gorbunov2024methods, vankov2024optimizing]. In this experiment, we keep $\bar{L}$ fixed and compute $L$ according to the rules established in Appendix D. The results are presented in Figure 2. For different values of $L$ and $\bar{L}$ there seems to be a trade-off between faster convergence to medium accuracy and slower convergence to very good accuracy for all three preconditioned methods.
Figure 2: Minimizing $\tfrac{1}{4}\|x\|^{4}$ using (2). The figure on the left corresponds to $\phi_{1}(x)=\cosh(\|x\|)-1$, the middle one to $\phi_{2}(x)=\exp(\|x\|)-\|x\|-1$ and the one on the right to $\phi_{3}(x)=-\|x\|-\ln(1-\|x\|)$.
@@ 4.2 Nonconvex phase retrieval
with $y_{i}\in\mathbb{R},a_{i}\in\mathbb{R}^{n}$. The data is generated as in [chen2020understanding]: $n=100$, $m=3000$ and $a_{i},z\sim\mathcal{N}(0,0.5)$, $x_{0}\sim\mathcal{N}(5,0.5)$ are generated element-wise with $z$ denoting the true underlying object. The measurements are generated as $y_{i}=(a_{i}^{\top}z)^{2}+n_{i}$ with $n_{i}\sim\mathcal{N}(0,4^{2})$.
We compare the algorithm (2) with the isotropic and separable reference functions generated by $\cosh-1$, denoted respectively by $\phi_{1}(x)=\cosh(\|x\|)-1$ and $\phi_{2}(x)=\sum_{i=1}^{n}\cosh(x_{i})-1$ against vanilla gradient descent, gradient clipping [zhang2019gradient] and [chen2020understanding, Algorithm 1] with the tuning described in [chen2020understanding, Section 7]. For the isotropic case of (2) we take $\gamma=5/3$ and $\lambda=1/100$, while for the separable one $\gamma=1/5$ and $\lambda=1/14$. The results are presented in Figure 3. We use as $f^{\star}$ the minimum value of the cost function among all algorithms. In this experiment the two versions of the algorithm proposed in this paper outperform the rest of the methods.
Moreover, we test the clipping and the isotropic algorithm over $100$ random instances of the problem. It can be seen that the isotropic algorithm outperforms the clipping method across the tests for this particular tuning. Note that the tuning for both of the methods is quite robust.
Figure 3: Nonconvex phase retrieval. $\phi_{1}$ corresponds to the isotropic reference function and $\phi_{2}$ to the separable one, both of which are generated by $\cosh(\cdot)-1$. The two figures on the left compare the algorithms for one instance of the problem. The figure on the right displays the results of gradient clipping and the isotropic version of (2) across 100 random instances.
@@ 5 Conclusion and Future Work
Our work paves the way for better understanding clipping and signed gradient methods from a majorization-minimization perspective. Possible interesting future work includes integrating momentum both in the convex and nonconvex regime and studying the stochastic setup.
[…]
@@ Appendix A Missing proofs of Section 1
To begin with, since $h$ is proper, lsc and convex, $h=h^{**}$. In light of [RoWe98, Theorem 11.8], we have that $\min h^{*}=-h^{**}(0)=-h(0)=0$, implying that $h^{*}\geq 0$. Moreover, from the same theorem, $\operatorname*{arg\,min}h=\{{h^{*}}^{\prime}(0)\}$ further implying that ${h^{*}}^{\prime}(0)=0$ and $h^{*}(0)=0$. Since $h$ is even, we have from [bauschke2017correction, Example 13.8] that $h^{*}=(h\circ|\cdot|)^{*}=h^{*}\circ|\cdot|$, which means that $h^{*}$ is also even. Therefore, through [bauschke2017correction, Proposition 11.7] we get that $h^{*}$ is increasing on $\mathbb{R}_{+}$.
Now, note that the function $g=h-\tfrac{\mu}{2}|\cdot|^{2}$ is proper, lsc and convex where $\mu$ is the strong convexity parameter of $h$. Moreover, it is even as the difference of two even functions. Therefore, from [bauschke2017correction, Proposition 11.7] $g$ is an increasing function on $\mathbb{R}_{+}$ and thus $g\circ\|\cdot\|$ is a convex function on $\mathbb{R}^{n}$. This implies that $h(\|x\|)-\tfrac{\mu}{2}\|x\|^{2}$ is convex and thus that $\phi=h\circ\|\cdot\|$ is strongly convex.
Again from [bauschke2017correction, Example 13.8], we have that $\phi^{*}=h^{*}\circ\|\cdot\|$ and to get the gradient of $\phi^{*}$ we can then utilize [bauschke2017correction, Corollary 16.72] to obtain $\nabla\phi^{*}(y)={h^{*}}^{\prime}(\|y\|)\overline{\operatorname{sgn}}(y)$.
Regarding the twice continuous differentiability of $\phi^{*}$, it follows from [rockafellar1977higher, p. 42] that $h^{*}\in\mathcal{C}^{2}(\mathbb{R}^{n})$ and the claimed result follows from [strichartz2000, Exercise 10.2.20], since $h^{*}$ is even. ∎
@@
It is straightforward that $h(x)=\tfrac{1}{2}x^{2}+\delta_{[-1,1]}(x)$, is a proper, lsc, strongly convex and even function with $h(0)=0$. Then, from [RoWe98, Theorem 11.23] we have that $h^{*}(y)=\inf_{x}\sigma_{[-1,1]}(x)+\tfrac{1}{2}(y-x)^{2}$, where $\sigma_{[-1,1]}$ is the support function of $[-1,1]$ and in light of [RoWe98, Exercise 11.27], ${h^{*}}^{\prime}(y)=\Pi_{[-1,1]}(y)$, where $\Pi_{[-1,1]}(y)=\min(1,\max(-1,y))$ is the projection on the closed convex set $[-1,1]$. Using Lemma 1.3 we thus obtain $\nabla\phi^{*}(y)=\min(1,\|y\|)\overline{\operatorname{sgn}}(y)$ and the algorithm becomes:
@@
by pulling the norm inside the $\min$.
@@ Appendix B Missing proofs of Section 2
Consider the following quantities, which are generalized conjugates as defined in [RoWe98, Chapter 11]:
@@
which is the claimed result for $\xi=(-f)^{\Phi}$. Now, assume moreover that $f$ is convex. In light of [RoWe98, Theorem 11.23] and [bauschke2017correction, Proposition 13.23] we have
$f^{*}=(\inf_{y\in\mathbb{R}^{n}}\bar{L}(L^{-1}\star\phi)(\cdot-y)+(-f)^{\Phi}(y))^{*}=\xi^{*}+(\bar{L}(L^{-1}\star\phi))^{*}=\xi^{*}+L^{-1}(\bar{L}\star\phi^{*})$
and as such the result follows for $\psi=\xi^{*}$ which is lsc and convex. ∎
@@
Consider any points $x,\bar{x}\in\mathbb{R}^{n}$. If $L(x-\bar{x}+L^{-1}\nabla\phi^{*}(\nabla f(\bar{x})))\notin\operatorname{dom}\phi$ then the bound holds trivially. Otherwise, from the descent lemma for $f$ we have:
@@
where we have used the fact that $\operatorname{rge}\nabla\phi^{*}\subseteq\operatorname{dom}\phi$ along with $\partial\phi(\nabla\phi^{*}(\nabla f(\bar{x})))=\{\nabla f(\bar{x})\}$. Therefore, the claimed result follows. ∎
@@
Let $H\coloneq\nabla^{2}f(x)$ and $Q\coloneq\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))\succ 0$, then
$H\preceq\bar{L}LQ^{-1}\Longleftrightarrow Q^{1/2}HQ^{1/2}\preceq\bar{L}LI\Longleftrightarrow\lambda_{\rm max}(Q^{1/2}HQ^{1/2})\leq\bar{L}L\Longleftrightarrow\lambda_{\rm max}(QH)\leq\bar{L}L$
where the first equivalence follows by [Horn_Johnson_2012, Theorem 7.7.2c] with $S=Q^{1/2}$ and the last equivalence follows from [Horn_Johnson_2012, Theorem 1.3.22]. The sufficient condition follows from Weyl’s inequality [Horn_Johnson_2012, Theorem 4.3.1].

### v2

@@ 3.2 Convex setting
Figure 3: Nonconvex phase retrieval. $\phi_{1}$ corresponds to the isotropic reference function and $\phi_{2}$ to the anisotropic one, both of which are generated by $\cosh(\cdot)-1$. The two figures on the left compare the algorithms for one instance of the problem. The figure on the right displays the results of gradient clipping and the isotropic version of (2) averaged across 100 random instances.
Figure 4: Simple NN training. (left) results for (2) with $\phi_{1}(x)=\cosh(\|x\|)-1$; (middle) $\phi_{2}(x)=-\|x\|-\ln(1-\|x\|)$; (right) gradient clipping method as presented in Example 1.7.
In the convex setting, Proposition 2.2 establishes a useful connection between anisotropic smoothness and the convexity of $f^{*}-L^{-1}(\bar{L}\star\phi^{*})$, i.e., the strong convexity of $f^{*}$ relative to $\bar{L}\star\phi^{*}$ with constant $L^{-1}$ in the Bregman sense. We utilize this connection in Propositions 3.5 and 3.6 in order to obtain standard sublinear convergence rates for the suboptimality gap in the isotropic case.
@@
Obtaining sublinear rates for the function values is not a straightforward task. To the best of our knowledge, there do not exist such guarantees for the full generality of the setting we consider in this paper. Proposition 3.5 is useful in that regard, since it allows us to show a $O(1/K)$ rate for the suboptimality gap in the isotropic case, as we show next.
@@
Let Assumption 3.1 hold, $f$ be convex and $\phi=h\circ\|\cdot\|$ with $h$ satisfying Assumption 1.1. For $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ the sequence of iterates generated by (2) with $\gamma=L^{-1}$ and $\lambda=\bar{L}^{-1}$, the following holds:
@@
We remark that $h{{}^{*}}^{\prime}(x)/x$ being a decreasing function on $\mathbb{R}_{+}$ is in fact a mild assumption that holds for all the kernel functions presented in Table 1. The result above strengthens the known results regarding the convergence of the method from (Maddison et al., 2021) and (Laude & Patrinos, 2025), while also answering the question posed in (Maddison et al., 2021, p. 17), regarding obtaining convergence guarantees for the suboptimality gap $f(x^{k})-f^{\star}$. We remark that although the obtained rate in (11) depends on the initial norm of the gradient, one can use the techniques from (Vankov et al., 2024) or (Gorbunov et al., 2024) to achieve a better complexity when specifying the reference function $\phi$. Nevertheless, such an endeavor is beyond the scope of this paper.
Although Theorem 3.6 provides a sublinear rate for the isotropic case, obtaining such guarantees for anisotropic reference functions is not straightforward: key in the proof of Theorem 3.6 is the fact that $\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k}))={h^{*}}^{\prime}(\|\bar{L}^{-1}\nabla f(x^{k})\|)\overline{\operatorname{sgn}}(\nabla f(x^{k}))$ and therefore the convex gradient inequality for $f$ can directly be utilized to show the Fejér monotonocity of $\{x^{k}\}_{k\in\mathbb{N}_{0}}$. Nevertheless, we are able to prove the sublinear convergence rate for the suboptimality gap for subhomogeneous (Azé & Penot, 1995, p. 708) reference functions using a different technique based on generalized conjugacy. More precisely, we utilize the characterization of anisotropically smooth functions as (generalized) envelopes and interpret the algorithm as a nonlinear proximal point method, generalizing thus the duality between gradient descent and proximal point in the Euclidean case (Laude, 2021, Theorem 3.8). Then, we combine the subhomogeneity of $\phi$ with the proof technique of (Doikov & Nesterov, 2020, Theorem 1) for inexact tensor methods and obtain the claimed rate. This result is captured in the following theorem.
@@
Let Assumption 3.1 hold, $f$ be convex and $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ be the sequence of iterates generated by (2) with $\gamma=L^{-1}$, $\lambda=\bar{L}^{-1}$ and assume that $\operatorname{dom}\phi=\mathbb{R}^{n}$. Moreover, let $\phi$ be $2$-subhomogeneous, i.e., such that $\phi(\theta x)\leq\theta^{2}\phi(x)$ for all $\theta\in[0,1]$. Then, for all $K\geq 1$
@@
where $\mathcal{D}_{0}=\sup\{\bar{L}(L^{-1}\star\phi)(x-x^{\star}):f(x)\leq f(x^{0})\}$ for $x^{\star}\in\operatorname*{arg\,min}f$.
In general the set $\mathcal{D}_{0}$ might be unbounded, except if $f$ has bounded level-sets, which is the case if $\operatorname*{arg\,min}f$ is bounded in light of (Bauschke et al., 2017b, Proposition 11.13). In theory, this dependence on the initial level-set can be eliminated by considering an averaging procedure akin to (Doikov & Nesterov, 2020, Algorithm 3), thus leading to a convergence rate in terms of some function of the initial distance to the solution. Nevertheless, such methods tend to underperform in practice compared to their more straightforward counterparts.
@@
The function $h(x):=\cosh(x)-1$ is $2$-subhomogeneous, i.e., the following inequality holds:
@@ 4 Experiments
In this section we present some simple experiments that display the behavior of the proposed method on problems beyond traditional Lipschitzian assumptions. The code for reproducing the experiments is publicly available [^1] https://github.com/JanQ/nonlinearly-preconditioned-gradient.
@@ 4.1 Norm to power
For the first part of our experiments we consider the toy example of minimizing $f(x)=\tfrac{1}{4}\|x\|^{4}$, with $x\in\mathbb{R}^{500}$, using different preconditioning schemes. We consider the reference functions $\cosh(\|x\|)-1$, $\exp(\|x\|)-\|x\|-1$ and $-\|x\|-\ln(1-\|x\|)$ and remind that the algorithm generated by the latter function is a tighter version of the algorithm proposed in (Gorbunov et al., 2024). In this experiment, we keep $\bar{L}$ fixed, compute $L$ according to the rules established in Appendix D and apply algorithm (2) with $\gamma=L^{-1}$ and $\lambda=\bar{L}^{-1}$. The results are presented in Figure 2. For different values of $L$ and $\bar{L}$ there seems to be a trade-off between faster convergence to medium accuracy and slower convergence to very good accuracy for all three preconditioned methods.
@@ 4.2 Nonconvex phase retrieval
with $y_{i}\in\mathbb{R},a_{i}\in\mathbb{R}^{n}$. The data is generated as in (Chen et al., 2023): $n=100$, $m=3000$ and $a_{i},z\sim\mathcal{N}(0,0.5)$, $x_{0}\sim\mathcal{N}(5,0.5)$ generated element-wise with $z$ denoting the true underlying object. The measurements are generated as $y_{i}=(a_{i}^{\top}z)^{2}+n_{i}$ with $n_{i}\sim\mathcal{N}(0,4^{2})$.
We compare the algorithm (2) with the isotropic and anisotropic reference functions generated by $\cosh-1$, denoted respectively by $\phi_{1}(x)=\cosh(\|x\|)-1$ and $\phi_{2}(x)=\sum_{i=1}^{n}\cosh(x_{i})-1$ against vanilla gradient descent, gradient clipping (Zhang et al., 2020b) and (Chen et al., 2023, Algorithm 1) with the tuning described in (Chen et al., 2023, Section 7). For the isotropic case of (2) we take $\gamma=5/3$ and $\lambda=1/100$, while for the anisotropic one $\gamma=1/5$ and $\lambda=1/14$. The results are presented in Figure 3. We use as $f^{\star}$ the minimum value of the cost function among all algorithms. In this experiment the two versions of the algorithm proposed in this paper outperform the rest of the methods.
Moreover, we test the clipping and the isotropic algorithm over $100$ random instances of the problem and plot the mean along with error bars representing one standard deviation on a logarithmic scale. It can be seen that the isotropic algorithm outperforms the clipping method across the tests for this particular tuning. Note that the tuning for both of the methods is quite robust.
4.3 Neural network training
In this experiment we consider training a simple four-layer fully connected network with layer dimensions $[28\times 28,128,64,32,32,10]$ and ReLU activation functions on a subset of the MNIST dataset (Deng, 2012), using the cross-entropy loss. We consider a subset ($m=600$) of the dataset in order to efficiently use full gradient updates.
We compare the methods generated by $\phi_{1}(x)=\cosh(\|x\|)-1$, $\phi_{2}(x)=-\|x\|-\ln(1-\|x\|)$ and the gradient clipping method (Zhang et al., 2020b), that can also be considered as an instance of (2) through Example 1.7, for various choices of the stepsizes and the clipping parameters. The results are presented in Figure 4. It can be seen that different combinations of $\gamma$ and $\lambda$ lead to different behaviors for the compared methods.
@@ 5 Conclusion and Future Work
Our work paves the way for better understanding clipping and signed gradient methods from a majorization-minimization perspective. Possible interesting future work includes integrating momentum both in the convex and nonconvex regime and studying the stochastic setup. Another interesting research direction is extending our convergence results for the suboptimality gap from the smooth to the additive nonsmooth setting where the nonsmooth term is handled similarly to (Laude & Patrinos, 2025). We believe that this extension is not straightforward and requires additional effort compared to the standard Euclidean setup of gradient descent.
Acknowledgements
Work supported by: the Research Foundation Flanders (FWO) research projects G081222N, G033822N, G0A0920N; Research Council KUL grant C14/24/103.
The authors thank Thomas Moellenhoff for helpful conversations.
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Missing proofs of Section 1
To begin with, since $h$ is proper, lsc and convex, $h=h^{**}$. In light of (Rockafellar & Wets, 1998, Theorem 11.8), we have that $\min h^{*}=-h^{**}(0)=-h(0)=0$, implying that $h^{*}\geq 0$. Moreover, from the same theorem, $\operatorname*{arg\,min}h=\{{h^{*}}^{\prime}(0)\}$ further implying that ${h^{*}}^{\prime}(0)=0$ and $h^{*}(0)=0$. Since $h$ is even, we have from (Bauschke et al., 2017b, Example 13.8) that $h^{*}=(h\circ|\cdot|)^{*}=h^{*}\circ|\cdot|$, which means that $h^{*}$ is also even. Therefore, through (Bauschke et al., 2017b, Proposition 11.7) we get that $h^{*}$ is increasing on $\mathbb{R}_{+}$.
Now, note that the function $g=h-\tfrac{\mu}{2}|\cdot|^{2}$ is proper, lsc and convex where $\mu$ is the strong convexity parameter of $h$. Moreover, it is even as the difference of two even functions. Therefore, from (Bauschke et al., 2017b, Proposition 11.7) $g$ is an increasing function on $\mathbb{R}_{+}$ and thus $g\circ\|\cdot\|$ is a convex function on $\mathbb{R}^{n}$. This implies that $h(\|x\|)-\tfrac{\mu}{2}\|x\|^{2}$ is convex and thus that $\phi=h\circ\|\cdot\|$ is strongly convex with the same strong convexity parameter.
Again from (Bauschke et al., 2017b, Example 13.8), we have that $\phi^{*}=h^{*}\circ\|\cdot\|$ and to get the gradient of $\phi^{*}$ we can then utilize (Bauschke et al., 2017b, Corollary 16.72) to obtain $\nabla\phi^{*}(y)={h^{*}}^{\prime}(\|y\|)\overline{\operatorname{sgn}}(y)$.
Regarding the twice continuous differentiability of $\phi^{*}$, it follows from (Rockafellar, 1977, p. 42) that $h^{*}\in\mathcal{C}^{2}(\mathbb{R}^{n})$ and the claimed result follows from (Strichartz, 2000, Exercise 10.2.20), since $h^{*}$ is even. ∎
@@
Proof.
It is straightforward that $h(x)=\tfrac{1}{2}x^{2}+\delta_{[-1,1]}(x)$, is a proper, lsc, strongly convex and even function with $h(0)=0$. Then, from (Rockafellar & Wets, 1998, Theorem 11.23) we have that $h^{*}(y)=\inf_{x}\sigma_{[-1,1]}(x)+\tfrac{1}{2}(y-x)^{2}$, where $\sigma_{[-1,1]}$ is the support function of $[-1,1]$ and in light of (Rockafellar & Wets, 1998, Exercise 11.27), ${h^{*}}^{\prime}(y)=\Pi_{[-1,1]}(y)$, where $\Pi_{[-1,1]}(y)=\min(1,\max(-1,y))$ is the projection on the closed convex set $[-1,1]$. Using Lemma 1.3 we thus obtain $\nabla\phi^{*}(y)=\min(1,\|y\|)\overline{\operatorname{sgn}}(y)$ and the algorithm becomes:
@@
by pulling the norm inside the $\min$. ∎
@@ Appendix B Missing proofs of Section 2
Consider the following quantities, which are generalized conjugates as defined in (Rockafellar & Wets, 1998, Chapter 11L):
@@
which is the claimed result for $\xi=(-f)^{\Phi}$.
We now show the convexity of $f^{*}-L^{-1}(\bar{L}\star\phi^{*})$. In light of (Rockafellar & Wets, 1998, Theorem 11.23) and (Bauschke et al., 2017b, Proposition 13.23) we have
$f^{*}=\big{(}\inf_{y\in\mathbb{R}^{n}}\bar{L}(L^{-1}\star\phi)(\cdot-y)+(-f)^{\Phi}(y)\big{)}^{*}=\xi^{*}+(\bar{L}(L^{-1}\star\phi))^{*}=\xi^{*}+L^{-1}(\bar{L}\star\phi^{*})$
and as such the result follows for $\psi=\xi^{*}$ which is lsc and convex, using moreover the fact that $\operatorname{dom}\phi^{*}=\mathbb{R}^{n}$ since $\phi$ is strongly convex. ∎
@@
Consider any points $x,\bar{x}\in\mathbb{R}^{n}$. If $L(x-\bar{x}+L^{-1}\nabla\phi^{*}(\nabla f(\bar{x})))\notin\operatorname{dom}\phi$ then the bound holds trivially. Otherwise, from the Euclidean descent lemma for $f$ we have:
@@
where we have used the fact that $\operatorname{rge}\nabla\phi^{*}\subseteq\operatorname{dom}\phi$ along with $\nabla f(\bar{x})\in\partial\phi(\nabla\phi^{*}(\nabla f(\bar{x})))$. Therefore, the claimed result follows. ∎
@@
Let $Q\coloneq\nabla^{2}f(x)$ and note that $H\coloneq\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))\succ 0$ from (Rockafellar, 1977, p. 42). Then
$Q\prec\bar{L}LH^{-1}\Longleftrightarrow H^{1/2}QH^{1/2}\prec\bar{L}LI\Longleftrightarrow\lambda_{\rm max}(H^{1/2}QH^{1/2})<\bar{L}L\Longleftrightarrow\lambda_{\rm max}(HQ)<\bar{L}L,$
where the first equivalence follows by (Horn & Johnson, 2012, Theorem 7.7.2c) with $S=H^{1/2}$. The last equivalence follows by noting that the (generally nonsymmetric) matrix $HQ$ is similar to the symmetric matrix $H^{1/2}QH^{1/2}$, by using (Horn & Johnson, 2012, Theorem 1.3.22) with $A=H^{1/2}Q$ and $B=H^{1/2}$, and noting that $H^{1/2}$ is nonsingular. The sufficient condition follows from Weyl’s inequality (Horn & Johnson, 2012, Theorem 4.3.1).